5. -4 points OSColPhys2016 18.3. P.024 My Notes Ask Your Teacher (a) Two point charges totaling 7.00 HC exert a repulsive force of 0.300 N on one another when separated by 0.352 m. What is the charge (in HC) on each? smallest charge HC largest charge b) What is the charge (in HC) on each if the force is attractive? smallest charge HC largest charge HC Additional Materials ReadingCan someone help me with this and provide a walkthrough for me as well? So I can better understand how to solve it on my own.
Answer #1

a)

let the two charges be q1 and q2

Given : Sum of charge = 7 uC

q1 + q2 = 7 x 10-6 C

q1 = (7 x 10-6) - q2                         eq-1

F = force of repusion between two charges = 0.300 N

r = distance between the charges = 0.352 m

using coulomb's law , electric force between two charges is given as

F = k q1 q2 /r2 ,

0.300 = (9 x 109) ((7 x 10-6) - q2 ) q2 /(0.352)2

q2 = 6.35 x 10-6 C = 6.35 uC

using eq-1

q1 = (7 x 10-6) - q2 = (7 x 10-6) - (6.35 x 10-6 ) = 0.65 x 10-6 C = 0.65 uC

smallest charge : 0.65 uC

largest charge : = 6.35 uC

b)

when force is attaractive

Given : Sum of charge = 7 uC

q1 - q2 = 7 x 10-6 C

q1 = (7 x 10-6) + q2                         eq-1

F = force of attaraction between two charges = 0.300 N

r = distance between the charges = 0.352 m

using coulomb's law , electric force between two charges is given as

F = k q1 q2 /r2 ,

0.300 = (9 x 109) ((7 x 10-6) + q2 ) q2 /(0.352)2

q2 = - 0.55 x 10-6 C = - 0.55 uC

using eq-1

q1 = (7 x 10-6) - q2 = (7 x 10-6) - (- 0.55 x 10-6 ) = 7.55 x 10-6 C = 7.55 uC

smallest charge : 0.55 uC

largest charge : = 7.55 uC